# On Supports for Graphs of Bounded Genus [^1] A preliminary version of these results appeared in EUROCOMB’23 [raman2023hypergraph]

### v3

@@ 1. $(i)$
    The vertices of $K$ are contained in the vertices $K_{0}$ in the arc $(a_{\ell},b_{j})$, and
@@ 2. $(ii)$
Suppose the lemma holds for all $(C^{\prime},\mathcal{K}^{\prime})$ with $\mathrm{cost}(C^{\prime},\mathcal{K}^{\prime})<N$. Given an instance with $\mathrm{cost}(C,\mathcal{K})=N$, by Lemma 24, there is a non-blocking chord $d=\{x,y\}$ joining two disjoint runs of some subgraph $K_{0}\in\mathcal{K}$.
@@
Hence, by the inductive hypothesis on the pair $(C_{\ell},\mathcal{K}_{\ell})$, there exists a set of non-intersecting chords $D_{\ell}$ such that each $K\in\mathcal{K}_{\ell}$ induces a connected subgraph of $C_{\ell}\cup D_{\ell}$. Similarly, there exists a set of non-intersecting chords $D_{r}$ such that each $K\in\mathcal{K}_{r}$, induces a connected subgraph of $C_{r}\cup D_{r}$. Let $D=D_{\ell}\cup D_{r}\cup d$. If there is a subgraph $K\in\mathcal{K}$ such that $K$ induces connected subgraphs on both $C_{r}\cup D_{r}$ and $C_{\ell}\cup D_{\ell}$ but $K$ is not connected on $C\cup D$, then $K\cap\{x,y\}=\emptyset$. It follows that $K$ is blocked by $d$; a contradiction that $d$ is a non-blocking chord. Hence, it $D$ is a set of non-intersecting chords such that each $K\in\mathcal{K}$ induces a connected subgraph of $(C\cup D)$.
@@ 6.1 Primal Support
Let $(G,\mathcal{H})$ be a graph system of genus $g$ such that each $H\in\mathcal{H}$ is a connected subgraph of $G$. Let $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ be any 2-coloring of $V(G)$ such that no vertex in $\mathbf{r}(V(G))$ is maximal and there are no adjacent twins $u,v\in\mathbf{r}(V(G))$.
@@
For each vertex $u\in\mathbf{r}(V(G))$, we do the following: Since $u$ is not maximal, by definition, there is an edge $e=\{u,v\}\in E(G)$ such that $\mathcal{H}_{e}=\mathcal{H}_{u}$. We chose one such edge $e$ incident to $u$, breaking ties arbitrarily. Since $\mathcal{H}_{e}\subseteq\mathcal{H}_{v}$, it follows that $\mathcal{H}_{u}\subseteq\mathcal{H}_{v}$. We orient this edge from $u$ to $v$. Further, if $v\in\mathbf{r}(V(G))$, then $\mathcal{H}_{u}\subset\mathcal{H}_{v}$ since there are no adjacent twins in $\mathbf{r}(V(G))$.
@@
Note that no red vertex is maximal in $G^{\prime}$ since each red vertex in $G^{\prime}$ has an outgoing edge incident on it. Further, there are no adjacent twins in $G^{\prime}$ among the red vertices as the edge contraction did not introduce new adjacencies between the red vertices. To see this note that the vertex obtained by contracting the edge $\{u,v\}$ is colored $\mathbf{b}$, and hence, the adjacency among the red vertices is unchanged. Since there are no adjacent red twins in $G$ it implies that there are no adjacent twins in $G^{\prime}$. Finally, note that $G^{\prime}$ has genus at most $g$ since contracting an edge in $G$ preserves the embedding on the same surface as $G$ does.
@@
Therefore, $(G^{\prime},\mathcal{H}^{\prime})$ is a cross-free graph system of genus $g$ such that the maximum depth of a maximal red vertex is at most k1k-1, and there are no adjacent red twins. By the induction hypothesis, $(G^{\prime},\mathcal{H}^{\prime})$ admits a primal support $Q$ of genus at most $g$, which is also a primal support for $(G,\mathcal{H})$ since the set of blue vertices remains unchanged for each subgraph in $\mathcal{H}$.
@@ 6.2 Dual Support
For each vertex $v\in V(G)$, the subgraphs containing $v$ trivially induce a connected subgraph in $Q^{*}$ since $0pt(v)=1$ for all $v\in V(G)$. Consider a special edge $e=\{u,v\}\in E(G)$. Since both end-points belong to a subgraph in $\mathcal{H}$, and $\mathcal{H}_{u}\cap\mathcal{H}_{v}=\emptyset$, it implies that $\mathcal{H}_{u}\not\subseteq\mathcal{H}_{v}$ and $\mathcal{H}_{v}\not\subseteq\mathcal{H}_{u}$. In the construction of $Q^{*}$ therefore, $e$ is not contracted. Further, $u$ and $v$ are contracted to vertices corresponding to the unique subgraph containing $u$ and $v$, respectively. Therefore, $Q^{*}$ satisfies the special edge property. ∎
@@ 6.3 Intersection Support
In this section, we show how we can construct an intersection support for a cross-free intersection system $(G,\mathcal{H},\mathcal{K})$. A vertex $v\in G$ s.t. $\mathcal{H}_{v}=\emptyset$, but $\mathcal{K}_{v}\neq\emptyset$ is called a $\mathcal{K}$-vertex, i.e., a vertex of $G$ that is contained in one or more subgraphs in $\mathcal{K}$ but none of the subgraphs in $\mathcal{H}$.
@@ 2. $(b)$
    where $|VC|^{*}$ denotes the size of an optimal vertex cover for the graph induced on the vertices of weight $1/2$.
@@ 7.2 Regions on an Oriented Surface
Figure 6: The subgraphs AA colored orange, $B$ colored green, and $C$ colored blue are touching. Observe that $A\setminus B$ is connected, while $B\setminus A$ induces two connected components. There is an input point on the boundary of the three regions. The regions are weakly non-piercing, but are not non-piercing.
@@
Moreover, by construction, any two regions $R_{1},R_{2}\in\mathcal{H}\cup\mathcal{K}$ intersect if and only if the corresponding subgraphs $R^{\prime}_{1},R^{\prime}_{2}\in\mathcal{H}^{\prime}\cup\mathcal{K}^{\prime}$ share a vertex in $G$. This implies the intersection hypergraph defined by $\mathcal{H}$ and $\mathcal{K}$, and that defined by $(G,\mathcal{H}^{\prime},\mathcal{K}^{\prime})$ are isomorphic.
@@ 7.3 Coloring Geometric Hypergraphs
By Theorem 14, there is a cross-free intersection system $(G,\mathcal{H}^{\prime},\mathcal{K}^{\prime})$ such that the intersection hypergraph defined by $\mathcal{H}$ and $\mathcal{K}$ is isomorphic to that defined $\mathcal{H}^{\prime}$ and $\mathcal{K}^{\prime}$. By Theorem 9, $(G,\mathcal{H}^{\prime},\mathcal{K}^{\prime})$ has an intersection support $\tilde{Q}$ of genus at most $g$. Now, $\chi(\tilde{Q})\leq\frac{7+\sqrt{1+24g}}{2}$ [diestel2005graph]. Since $\tilde{Q}$ is a support, for each $K\in\mathcal{K^{\prime}}$, there is an edge between some two subgraphs $H,H^{\prime}\in{\mathcal{H}^{\prime}_{K}}$. Therefore, no hyperedge $\mathcal{H}^{\prime}_{K}$} is monochromatic. ∎

### v5

@@ 1. $(i)$
    The vertices of $K$ are contained in the vertices of $K_{0}$ in the arc $(a_{\ell},b_{j})$, and
@@ 2. $(ii)$
Suppose the lemma holds for all $(C^{\prime},\mathcal{K}^{\prime})$ with $\mathrm{cost}(C^{\prime},\mathcal{K}^{\prime})<N$. Now, consider an instance with an embedded cycle $C$ and a set $\mathcal{K}$ of abababab-free subgraphs of $C$ such that $\mathrm{cost}(C,\mathcal{K})=N$. By Lemma 24, there is a non-blocking chord $d=\{x,y\}$ joining two disjoint runs of some subgraph $K_{0}\in\mathcal{K}$.
@@
Hence, by the inductive hypothesis on the pair $(C_{\ell},\mathcal{K}_{\ell})$, there exists a set of non-intersecting chords $D_{\ell}$ such that each $K\in\mathcal{K}_{\ell}$ induces a connected subgraph of $C_{\ell}\cup D_{\ell}$. Similarly, there exists a set of non-intersecting chords $D_{r}$ such that each $K\in\mathcal{K}_{r}$, induces a connected subgraph of $C_{r}\cup D_{r}$. Let $D=D_{\ell}\cup D_{r}\cup d$. If there is a subgraph $K\in\mathcal{K}$ such that $K$ induces connected subgraphs on both $C_{r}\cup D_{r}$ and $C_{\ell}\cup D_{\ell}$ but $K$ is not connected on $C\cup D$, then $K\cap\{x,y\}=\emptyset$. It follows that $K$ is blocked by $d$; a contradiction that $d$ is a non-blocking chord. Hence, $D$ is a set of non-intersecting chords such that each $K\in\mathcal{K}$ induces a connected subgraph of $(C\cup D)$.
@@ 6.1 Primal Support
Let $(G,\mathcal{H})$ be a graph system of genus $g$ such that each $H\in\mathcal{H}$ is a connected subgraph of $G$. Let $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ be any 2-coloring of $V(G)$ such that no vertex in $\mathbf{r}(V(G))$ is maximal and there are no adjacent twins $u,v\in\mathbf{r}(V(G))$. Then there is a primal support of genus at most $g$.
@@
For each vertex $u\in\mathbf{r}(V(G))$, we do the following: Since $u$ is not maximal, by definition, there is an edge $e=\{u,v\}\in E(G)$ such that $\mathcal{H}_{e}=\mathcal{H}_{u}$. We choose one such edge $e$ incident to $u$, breaking ties arbitrarily. Since $\mathcal{H}_{e}\subseteq\mathcal{H}_{v}$, it follows that $\mathcal{H}_{u}\subseteq\mathcal{H}_{v}$. We orient this edge from $u$ to $v$. Further, if $v\in\mathbf{r}(V(G))$, then $\mathcal{H}_{u}\subset\mathcal{H}_{v}$ since there are no adjacent twins in $\mathbf{r}(V(G))$.
@@
Note that no red vertex is maximal in $G^{\prime}$ since each red vertex in $G^{\prime}$ has an outgoing edge incident on it. Further, there are no adjacent twins in $G^{\prime}$ among the red vertices as the edge contraction did not introduce new adjacencies between the red vertices. To see this note that the vertex obtained by contracting the edge $\{u,v\}$ is colored $\mathbf{b}$, and hence, the adjacency among the red vertices is unchanged. Since there are no adjacent red twins in $G$ it implies that there are no adjacent twins in $G^{\prime}$. Finally, note that $G^{\prime}$ has genus at most $g$ since contracting an edge in $G$ preserves the embedding of the resulting graph on the same surface as that of $G$.
@@
Therefore, $(G^{\prime},\mathcal{H}^{\prime})$ is a cross-free graph system of genus $g$ such that the maximum depth of a maximal red vertex is at most k1k-1, and there are no adjacent red twins. By the induction hypothesis, $(G^{\prime},\mathcal{H}^{\prime})$ admits a primal support $Q$ of genus at most $g$, which is also a primal support for $(G,\mathcal{H})$ since the set of blue vertices remains unchanged for each subgraph in $\mathcal{H}$. ∎
@@ 6.2 Dual Support
For each vertex $v\in V(G)$, the subgraphs containing $v$ trivially induce a connected subgraph in $Q^{*}$ since $0pt(v)\leq 1$ for all $v\in V(G)$. Consider a special edge $e=\{u,v\}\in E(G)$. Since each of the two endpoints belong to a subgraph in $\mathcal{H}$, and $\mathcal{H}_{u}\cap\mathcal{H}_{v}=\emptyset$, it implies that $\mathcal{H}_{u}\not\subseteq\mathcal{H}_{v}$ and $\mathcal{H}_{v}\not\subseteq\mathcal{H}_{u}$. In the construction of $Q^{*}$ therefore, $e$ is not contracted. Further, $u$ and $v$ are contracted to vertices corresponding to the unique subgraph containing $u$ and $v$, respectively. Therefore, $Q^{*}$ satisfies the special edge property. ∎
@@ 6.3 Intersection Support
In this section, we show how we can construct an intersection support for a cross-free intersection system $(G,\mathcal{H},\mathcal{K})$. A vertex $v\in V(G)$ s.t. $\mathcal{H}_{v}=\emptyset$, but $\mathcal{K}_{v}\neq\emptyset$ is called a $\mathcal{K}$-vertex, i.e., a vertex of $G$ that is contained in one or more subgraphs in $\mathcal{K}$ but none of the subgraphs in $\mathcal{H}$.
@@ 2. $(b)$
    where $|VC^{*}|$ denotes the size of an optimal vertex cover for the graph induced on the vertices of weight $1/2$.
@@ 7.2 Regions on an Oriented Surface
Figure 6: The regions AA colored orange, $B$ colored green, and $C$ colored blue are touching. Observe that $A\setminus B$ is connected, while $B\setminus A$ induces two connected components. There is an input point on the boundary of the three regions. The regions are weakly non-piercing, but are not non-piercing.
@@
Moreover, by construction, any two regions $R_{1},R_{2}\in\mathcal{H}\cup\mathcal{K}$ intersect if and only if the corresponding subgraphs $R^{\prime}_{1},R^{\prime}_{2}\in\mathcal{H}^{\prime}\cup\mathcal{K}^{\prime}$ share a vertex in $G$. This implies the intersection hypergraph defined by $\mathcal{H}$ and $\mathcal{K}$, and that defined by $(G,\mathcal{H}^{\prime},\mathcal{K}^{\prime})$ are isomorphic. ∎
@@ 7.3 Coloring Geometric Hypergraphs
By Theorem 14, there is a cross-free intersection system $(G,\mathcal{H}^{\prime},\mathcal{K}^{\prime})$ such that the intersection hypergraph defined by $\mathcal{H}$ and $\mathcal{K}$ is isomorphic to that defined by $\mathcal{H}^{\prime}$ and $\mathcal{K}^{\prime}$. By Theorem 9, $(G,\mathcal{H}^{\prime},\mathcal{K}^{\prime})$ has an intersection support $\tilde{Q}$ of genus at most $g$. Now, $\chi(\tilde{Q})\leq\frac{7+\sqrt{1+24g}}{2}$ [diestel2005graph]. Since $\tilde{Q}$ is a support, for each $K\in\mathcal{K^{\prime}}$, there is an edge between some two subgraphs $H,H^{\prime}\in{\mathcal{H}^{\prime}_{K}}$. Therefore, no hyperedge $\mathcal{H}^{\prime}_{K}$} is monochromatic. ∎